# Quantifying uncertainty in climate projections with conformal ensembles

### v1

@@ Abstract
We introduce conformal ensembling, a new approach to uncertainty quantification in climate projections based on conformal inference. Unlike traditional methods, conformal ensembling seamlessly integrates climate models and observational data across a range of scales to generate statistically rigorous, easy-to-interpret uncertainty estimates. It can be applied to any climatic variable using any ensemble analysis method and outperforms existing inter-model variability methods in uncertainty quantification across all time horizons and most spatial locations under SSP2-4.5. Conformal ensembling is also computationally efficient, requires minimal assumptions, and is highly robust to the conformity measure. Experiments show that it is effective when conditioning future projections on historical reanalysis data compared with standard ensemble averaging approaches, yielding more physically consistent projections.
@@ 1 Introduction
Climate modeling centers actively develop new climate models, and improve old ones, to improve their capacity to emulate the global climate system (O’Neill et al., 2016). Due to variations between models, they can produce a wide range of plausible projections, even given the same initial conditions and forcings (Knutti et al., 2010; Flato et al., 2014). Current practice is to consider a multi-model ensemble that contains output from many different climate models run under the same time-evolving external forcings. A multi-model ensemble helps ensure that estimates of the climate response are robust across different model structures and allows us to quantify projection uncertainty through the inter-model variability (IMV) (Lambert and Boer, 2001; Gleckler, Taylor and Doutriaux, 2008; Knutti et al., 2010; Harris, Li and Sriver, 2023). The multi-model ensemble is then aggregated into a single projection, called an analysis, and paired with some notion of uncertainty, such as a 5%-95% model spread (Masson-Delmotte et al., 2021). This type of analysis is ubiquitous in climate science (Masson-Delmotte et al., 2021) and is used extensively to estimate impacts and risks associated with climate change (Jay et al., 2023; Masson-Delmotte et al., 2021).
@@
The degree to which large ensembles represent projection uncertainty is a topic of significant recent interest (Jones, 2000; Christensen et al., 2008; White and Toumi, 2013; Qian et al., 2016; Tett et al., 2022). Key sources of uncertainty include differences in model structures, future greenhouse scenarios, and natural variability within the climate system (Hawkins and Sutton, 2009). Uncertainties can also arise from unresolved process parameterizations, parameter value uncertainty, and even the data used for model calibration Tett et al. (2022). These uncertainties tend to increase as climate models become more complex, finer scale, and computationally expensive, creating challenges for projecting future regional variability and extremes (Shaw et al., 2024). This has led to a somewhat paradoxical situation where CMIP6 models show increased uncertainty in the future compared to CMIP5 models, with even more pronounced differences at finer scales.
@@
Recent works have proposed various methods for reducing projection uncertainty by direct conditioning on observational data. One common approach is to use bias correction methods (Piani, Haerter and Coppola, 2010; Teutschbein and Seibert, 2012; Vrac and Friederichs, 2015; Maraun et al., 2017) that shift and scale the ensemble to match the quantiles, or other functionals, of historical observational data. Bias correction methods can effectively reduce uncertainty in the mean state for the historical and current climate. Still, they can often result in overly confident projections and are sensitive to the method and target observational data sets (Lafferty and Sriver, 2023). Another approach includes Bayesian methods (Tebaldi et al., 2004; Smith et al., 2009; Bhat et al., 2011; Rougier, Goldstein and House, 2013; Qian et al., 2016; Sansom, Stephenson and Bracegirdle, 2017; Bowman et al., 2018) that taxonomize and quantify the different sources of variability in climate models (e.g., model uncertainty, model inadequacy, and natural variability). They can then learn a complex posterior distribution over the climatology conditional on observational data. Bayesian approaches offer a rigorous uncertainty quantification scheme but typically at an inordinately high computational cost without a surrogate model.
@@
We propose an alternative approach to uncertainty quantification based on conformal inference (Vovk, Gammerman and Shafer, 2005; Shafer and Vovk, 2008). Conformal inference is a general predictive inference methodology for imbuing any black box model with statistically valid prediction regions, even in finite samples (Lei et al., 2018). We re-cast the problem of climate projection as a prediction-like problem (Harris, Li and Sriver, 2023), whereby an “ensemble analysis function”, such as a deep neural network, is trained to predict observational data from a multi-model ensemble. The predictions of the analysis function constitute a synthesis of the inputted multi-model ensemble to a single projection that has been bias-corrected and downscaled against observational data. By reserving a small hold-out set of historical data, we can estimate the out-of-sample residual distribution of the analysis function and use this to form prediction sets for any future multi-model ensembles run through the same analysis function (Lei et al., 2018). We show that these prediction sets, represented as conformal ensembles, can be used in place of a multi-modal ensemble as a valid measure of climate projection uncertainty with significantly improved UQ skill.
@@
Conformal ensembles offer a number of advantages over traditional large ensemble approaches (Stocker, 2014) and Bayesian approaches that use the multi-model ensembles directly. Instead of attempting to estimate each model component’s uncertainty or the different variability sources, our framework shortcuts directly to the projection uncertainty. This relatively simple approach seamlessly integrates models and observational data across a wide range of scales into a rigorous and easy-to-understand measure of uncertainty. Compared to traditional inter-model variability based approaches, conformal ensembles are not subject to the structural differences across models and simplifying assumptions that lead to under-confidence. Compared to Bayesian approaches, they are computationally efficient, do not require subjective priors, and are universally applicable to any model analysis function. In fact, conformal ensembles do not require assumptions about the correctness of the models, the relationship between models, or the relationship between models and data. Our only requirement is that the projection residuals are exchangeable (Lei and Wasserman, 2014) over time, an assumption that we demonstrate approximately holds for many analysis functions across many climate variables (Section 4).
@@ 2.2 Reanalysis data
Instead of proper observational data, which is sparsely observed in time and space, we will use quasi-observational data, which has complete spatiotemporal coverage. Complete spatiotemporal coverage is necessary for our method to produce projection ensembles for the entire globe. Specifically, we use reanalysis data from the ERA5 reanalysis product (Hersbach et al., 2020). Reanalysis data optimally combines observational data from a wide variety of observational sources with weather models to produce high-resolution gridded estimates. ERA5 uses Integrated Forecasting System (IFS) Cy41r2 and spans January 1940 through March 2024. We will only consider TAS, TMAX, and PR fields on single pressure levels to match the CMIP6 runs (Section 2.1). However, unlike the CMIP6 runs, ERA5 interpolates on a much finer grid (30km) that we will not downscale to match the models. A benefit of our conformal approach is that the conformal ensemble will match the resolution of the target process (reanalysis fields), which results in an automatic downscaling effect.
@@ 3.2 Scoring multi-model ensemble forecasts
that includes all points, $r\in\mathbb{R}^{q}$ with a depth value greater than $\tau$. Central regions are key to how we implement the general ICP algorithm (Algorithm 1) for climate model projections (Algorithm 2). We will, essentially, just take $\tau$ to be the $\lceil(n_{1}+1)\alpha\rceil+1)$ smallest depth value and use the $D_{\tau}$ central region, translated by $\hat{f}_{\theta}$, as our prediction region $C_{n}(Z)$ (Algorithm 1 step 5). This procedure mimics the ordinary conformal prediction procedure for regression (Lei et al., 2018), but for high-dimensional multivariate targets using data depth for ranking.
@@ 4 Experiments
We conduct “perfect model” experiments to assess the coverage and UQ skill of the proposed conformal approach across a range of climate variables, ensemble analysis functions, and depth functions (Knutti et al., 2017; Harris, Li and Sriver, 2023). In each experiment we have an ensemble of $m$ climate model runs (TAS: $m=31$, TMAX: $m=20$, PR: $m=30$), which we split into a train set (1940 - 2006), calibration set (2007 - 2024), a test set (2024-2099). Iterating over all $m$ model runs, we treat one run as the target process, or stand-in, for observational data and use the remaining $m-1$ model runs as our multi-model ensemble. We train six different ensemble analysis functions and construct, for each, three different conformal ensembles based on three different depth functions on train and calibration sets respectively (Section 2.1). We compare the UQ skill of each combination against the inter-model variability (IMV) and bias-corrected inter-model variability (IMV (BC)) (Vrac and Friederichs, 2015) on the test set. By repeating this procedure for all $m$ model runs, we can assess conformal uncertainty quantification across a wide range of plausible climate trajectories.
@@
The three depth functions we consider include the $\ell_{\infty}$-Depth based on the $\ell_{\infty}$-norm ($\ell_{\infty}$), the Integrated Tukey depth (Tukey), and the (inverse) $\ell_{\infty}$ norm (norm), which is a non-depth based scoring rule. The $\ell_{\infty}$-Depth of a vector $r$ with respect to distribution $P$ is defined as $D_{\infty}(r,P)=1/(1+\mathbb{E}[|r-R|_{\infty}])$ where $R\sim P$. The Integrated Tukey depth averages the univariate Tukey depth $D_{\text{Tukey}}(r,P)=\mathbb{E}[\min\{\hat{F}_{x}(r(x),1-\hat{F}_{x}(r(x))\}]$ over each element of $r$. Here $\hat{F}_{x}$ denotes the empirical CDF of the residuals at location $x$. Finally, the $\ell_{\infty}$-norm simply computes $D(x)=|x|_{\infty}$. We choose $\ell_{\infty}$ norm based depths and scoring rules, instead of the usual $\ell_{1}$ or $\ell_{2}$ norm based measures because $\ell_{\infty}$ was found to be more robust to changes in the distribution of the ensembles.
@@ 4.1 Coverage and sharpness
Table 1: Out of sample (May 2024 - Dec 2099) empirical coverage, and average width, of each model analysis function (EA, WEA, Delta, LM, GP, CNN) on TAS, TMAX, and PR and a white noise (WN) control run averaged over all perfect model experiments. We included a white noise control run to assess the empirical coverage of conformal inference when all assumptions are met. Climate variable results show that conformal prediction controls the out-of-sample coverage near the nominal level, with a minor few exceptions. Proper depth functions ($\ell_{\infty}$-Depth and Tukey depth) show nominal to slightly conservative coverage.
@@ 4.2 Uncertainty Quantification
Table 2: Sliced Wasserstein distance between the projected distribution and the target process for each model analysis function and UQ method over the projection horizon (May 2024 - Dec 2099). Each number represents the average SW distance across all perfect model experiments and the number in parenthesis is one standard error. Regardless of the analysis function, depth function, or climatic variable, conformal ensembles show lower projection error compared to the IMV and IMV(BC).
@@
A major concern with bias correcting methods and other alternative approaches to UQ is the problem of overconfidence. We pose that the improvements seen in Table 2, using true depth functions ($\ell_{\infty}$ and Tukey), are not likely caused by overconfidence of the CE ensemble because, as was shown in Table 1, conformal methods have well controlled empirical coverage. The CE ensemble is drawn uniformly from the $1-\alpha$ confidence region (Section 3.3), so it is not simply concentrated near the median. In fact, the results in Table 1 suggest that the conformal prediction sets may even be slightly underconfident because the empirical sizes are generally greater than the nominal level. CE (Norm) shows slightly improved metrics compared with $\ell_{\infty}$-Depth and Tukey depth, but due to its consistent under-coverage towards the latter part of the prediction interval (2070-2100), this may be due to overconfidence.
@@ 4.3 Uncertainty quantification over time
As in Section 4.1, we break the aggregate UQ skill results in Table 2 down into decadal blocks to assess UQ skill over time. Because conformal methods can be susceptible to distribution shift, this will allow us to identify if there are any periods where CE performs worse than IMV or IMV(BC). We then further decompose our temporal results into monthly units, where SW is computed using only prediction sets and observations within a given month to assess seasonal variability. For CE, we use the $\ell_{\infty}$-Depth due to its overall more stable coverage (Figure 1) and comparable UQ skill compared to the other depths.
@@
The top row of Figure 2 shows the SW distance between the projection and the target over a given decade using each UQ method (IMV, IMV(BC), and CE) for each ensemble analysis function averaged over all model experiments. Although SW metrics tend to increase over time, particularly for TAS and TMAX projections, CE is consistently below IMV and IMV(BC) for all models and all time periods. Thus, even for far future projections (2070-2100), the climatologies produced by CE are much closer to the climatologies of the target process than if IMV had been used. One notable exception is LM using CE, which sits higher than all other models using CE in TAS and TMAX. This was likely caused by instabilities in the fitting process of LM, resulting in overly high or low projections at a small handful of spatial locations.
@@ 4.4 Marginal uncertainty quantification over space
Table 2 shows that, regardless of the ensemble analysis function or climate variable, swapping out IMV or IMV(BC) for a conformal ensemble (CE) uniformly decreases the approximation error of the projected climatology to the target processes climatology. Figure 2 confirms that this result holds for each decade in the projection interval (May 2024 - Dec 2099). Because these results are based on the Sliced Wasserstein (SW) distance, they measure how close the projection’s joint distribution is to the target process’s joint distribution. They do not tell us how close their distributions are marginally at each grid location, i.e., if there are any spatially varied differences.

### v2

@@ Abstract
Large climate model ensembles are the primary tool for robustly projecting future climate states and quantifying projection uncertainty. Despite significant advancements in climate modeling over the past few decades, overall projection certainty has not commensurately decreased with steadily improving model skill. We introduce conformal ensembling, a new approach to uncertainty quantification in climate projections based on conformal inference to reduce projection uncertainty. Unlike traditional methods, conformal ensembling seamlessly integrates climate model ensembles and observational data across a range of scales to generate statistically rigorous, easy-to-interpret uncertainty estimates. It can be applied to any climatic variable using any ensemble analysis method and outperforms existing inter-model variability methods in uncertainty quantification across all time horizons and most spatial locations under SSP2-4.5. Conformal ensembling is also computationally efficient, requires minimal assumptions, and is highly robust to the conformity measure. Experiments show that it is effective when conditioning future projections on historical reanalysis data compared with standard ensemble averaging approaches, yielding more physically consistent projections.
@@ 1 Introduction
Climate modeling centers actively develop new climate models, and improve old ones, to improve their capacity to emulate the global climate system (O’Neill et al., 2016). Due to variations between models, they can produce a wide range of plausible projections, even given the same initial conditions and forcings (Knutti et al., 2010; Flato et al., 2014). Current practice is to consider a multi-model ensemble that contains output from many different climate models run under the same time-evolving external forcings. A multi-model ensemble helps ensure that estimates of the climate response are robust across different model structures and allows us to quantify projection uncertainty through the inter-model variability (IMV) (Lambert and Boer, 2001; Gleckler, Taylor and Doutriaux, 2008; Knutti et al., 2010; Stocker, 2014). The multi-model ensemble is then aggregated into a single projection, called an analysis, and paired with some notion of uncertainty, such as a 5%-95% model spread (Masson-Delmotte et al., 2021). This type of analysis is ubiquitous in climate science (Masson-Delmotte et al., 2021) and is used extensively to estimate impacts and risks associated with climate change (Jay et al., 2023; Masson-Delmotte et al., 2021).
@@
The degree to which large ensembles represent projection uncertainty is a topic of significant recent interest (Jones, 2000; Christensen et al., 2008; White and Toumi, 2013; Qian et al., 2016; Tett et al., 2022). Key sources of uncertainty in climate projections include differences in model structure, model inadequacy, and the inherent internal variability within a model (Hawkins and Sutton, 2009; Tett et al., 2022). Uncertainties can also arise from unresolved process parameterizations, parameter value uncertainty, and even the data used for model calibration (Tett et al., 2022). These uncertainties tend to increase as climate models become more complex, resolved on finer scales, and more computationally expensive, creating challenges for projecting future regional variability and extremes (Shaw et al., 2024). This has even led to a somewhat paradoxical situation where newer models, drawn from the Coupled Model Intercomparison Project 6 (CMIP6) (Eyring et al., 2016), can show increased uncertainty compared to older CMIP5 models, particularly at regional scales.
@@
Recent works have proposed various methods for reducing projection uncertainty by direct conditioning on observational data. One common approach is to use bias correction methods (Piani, Haerter and Coppola, 2010; Teutschbein and Seibert, 2012; Vrac and Friederichs, 2015; Maraun et al., 2017) that shift and scale the ensemble to match the quantiles, or other functionals, of historical observational data. Bias correction methods are effective in reducing uncertainty in the mean states of the historical and current climate. However, they can often result in overly confident projections and are sensitive to the method and target observational data sets (Lafferty and Sriver, 2023). An alternative approach are Bayesian hierarchical models (Tebaldi et al., 2004; Smith et al., 2009; Bhat et al., 2011; Rougier, Goldstein and House, 2013; Qian et al., 2016; Sansom, Stephenson and Bracegirdle, 2017; Bowman et al., 2018) that taxonomize and quantify the different sources of variability in climate models (e.g., model uncertainty, model inadequacy, and natural variability). Bayesian methods learn a posterior distribution of local or global climatologies conditional on model runs and observational data. This process yields rigorous uncertainty quantification via posterior sampling, but typically at a high computational cost, using complex models, with subjectively chosen priors.
@@
We propose an alternative approach to uncertainty quantification based on conformal inference (Vovk, Gammerman and Shafer, 2005; Shafer and Vovk, 2008). Conformal inference is a general predictive inference methodology for imbuing any black box model with statistically valid prediction regions, even in finite samples (Lei et al., 2018). We re-cast the problem of climate projection as a prediction-like problem (Harris, Li and Sriver, 2023), whereby an “ensemble analysis function”, such as a deep neural network, is trained to predict observational data from a multi-model ensemble. The predictions of the analysis function constitute a synthesis of the inputted multi-model ensemble into a single projection that has been bias-corrected and downscaled against observational data. By reserving a small hold-out set of historical data, we can estimate the out-of-sample residual distribution of the analysis function and use this to form prediction sets for any future multi-model ensembles run through the same analysis function (Lei et al., 2018). We show that these prediction sets, represented as conformal ensembles, can be used in place of a multi-modal ensemble as a valid measure of climate projection uncertainty with significantly improved UQ skill.
@@
Conformal ensembles offer a number of advantages over traditional large ensemble and Bayesian approaches because they do not attempt to estimate each source of uncertainty, but rather directly shortcut to the final projection uncertainty. This relatively simple scheme, however, allows us to seamlessly integrate models runs and observational data over a wide range of spatial scales, using standard statistical learning approaches, into rigorous and easy-to-understanding measures of uncertainty. Compared to traditional inter-model variability based approaches, conformal ensembles are not subject to the structural differences across models and simplifying assumptions that lead to under-confidence. Compared to Bayesian approaches, they are computationally efficient, do not require subjective priors, and are universally applicable to any model analysis function. In fact, conformal ensembling does not require strong assumptions about the correctness of the models, the relationship between models, or the relationship between models and data. Our only requirement is that the projection residuals are exchangeable (Lei and Wasserman, 2014) over time, or equivalently that the relationship between the models and the observations is relatively stationary over time. We demonstrate that this assumption approximately holds for many analysis functions across many climate variables (Section 4).
@@ 2.2 Reanalysis data
Instead of proper observational data, which is sparsely observed in time and space, we will use quasi-observational data, which has complete spatiotemporal coverage. Complete spatiotemporal coverage is necessary for our method to produce projection ensembles for the entire globe. Specifically, we use reanalysis data from the ERA5 reanalysis product (Hersbach et al., 2020). Reanalysis data optimally combines observational data from a wide variety of observational sources with weather models to produce high-resolution gridded estimates. ERA5 uses Integrated Forecasting System (IFS) Cy41r2 and spans January 1940 through March 2024. We will only consider TAS, TMAX, and PR fields on single pressure levels to match the CMIP6 runs (Section 2.1). However, unlike the CMIP6 runs, ERA5 interpolates on a much finer grid (30km) that do not necessarily need to downscale to match the model’s resolution. A benefit of our conformal approach is that the conformal ensemble will match the resolution of the target process (reanalysis fields), which results in an automatic downscaling effect.
@@ 3.2 Scoring multi-model ensemble forecasts
that includes all points, $r\in\mathbb{R}^{q}$ with a depth value greater than $\tau$. Central regions are key to how we implement the general ICP algorithm (Algorithm 1) for climate model projections (Algorithm 2). We will, essentially, just take $\tau$ to be the $\lceil(n_{1}+1)\alpha\rceil+1)$ smallest depth value and use the $D_{\tau}$ central region, translated by $\hat{f}_{\theta}$, as our prediction region $C_{n}(Z)$ (Algorithm 1 step 5). This procedure is the ordinary conformal prediction procedure for regression (Lei et al., 2018), but for high-dimensional multivariate targets using data depth for ranking.
@@ 4 Experiments
We conduct “perfect model” experiments to assess the coverage and UQ skill of the proposed conformal approach across a range of climate variables, ensemble analysis functions, and depth functions (Knutti et al., 2017; Harris, Li and Sriver, 2023). In each experiment we have an ensemble of $m$ climate model runs (TAS: $m=31$, TMAX: $m=20$, PR: $m=30$), which we split into a train set (Jan. 1940 - Jun. 2007), calibration set (Jul. 2007 - Mar. 2024), a test set (Apr. 2024 - Dec. 2099). Iterating over all $m$ model runs, we treat one run as the target process, or stand-in, for observational data and use the remaining $m-1$ model runs as our multi-model ensemble. We train six different ensemble analysis functions and construct, for each, three different conformal ensembles based on three different depth functions on train and calibration sets respectively (Section 2.1). We compare the UQ skill of each combination against the inter-model variability (IMV) and bias-corrected inter-model variability (IMV (BC)) (Vrac and Friederichs, 2015) on the test set. By repeating this procedure for all $m$ model runs, we can assess conformal uncertainty quantification across a wide range of plausible climate trajectories.
@@
The three depth functions we consider include the $\ell_{\infty}$-Depth based on the $\ell_{\infty}$-norm, the Integrated Tukey depth (Tukey), and the (inverse) $\ell_{\infty}$ norm (norm), which is a non-depth based scoring rule. The $\ell_{\infty}$-Depth of a vector $r$ with respect to distribution $P$ is defined as $D_{\infty}(r,P)=1/(1+\mathbb{E}[||r-R||_{\infty}])$ where $R\sim P$. The Integrated Tukey depth averages the univariate Tukey depth $D_{\text{Tukey}}(r,P)=\mathbb{E}[\min\{\hat{F}_{x}(r(x),1-\hat{F}_{x}(r(x))\}]$ over each element of $r$. Here $\hat{F}_{x}$ denotes the empirical CDF of the residuals at location $x$. Finally, the $\ell_{\infty}$-norm simply computes $D(x)=1/(1+||x||_{\infty})$. We choose $\ell_{\infty}$-norm based depths and scoring rules, instead of the usual $\ell_{1}$ or $\ell_{2}$ norm based measures because $\ell_{\infty}$ was found to be more robust to changes in the distribution of the ensembles.
@@ 4.1 Coverage and sharpness
Table 1: Out of sample (Apr. 2024 - Dec 2099) empirical coverage, and average width, of each model analysis function (EA, WEA, Delta, LM, GP, CNN) on TAS, TMAX, and PR and a white noise (WN) control run averaged over all perfect model experiments. We included a white noise control run to assess the empirical coverage of conformal inference when all assumptions are met. Climate variable results show that conformal prediction controls the out-of-sample coverage near the nominal level, with a minor few exceptions. Proper depth functions ($\ell_{\infty}$-Depth and Tukey depth) show nominal to slightly conservative coverage.
@@ 4.2 Uncertainty Quantification
Table 2: Sliced Wasserstein distance between the projected distribution and the target process for each model analysis function and UQ method over the projection horizon (Apr. 2024 - Dec 2099). Each number represents the average SW distance across all perfect model experiments and the number in parenthesis is one standard error. Regardless of the analysis function, depth function, or climatic variable, conformal ensembles show lower projection error compared to the IMV and IMV(BC).
@@
A major concern with bias correcting methods and other alternative approaches to UQ is the problem of overconfidence. We pose that the improvements seen in Table 2, using true depth functions ($\ell_{\infty}$-Depth and Tukey), are not likely caused by overconfidence of the CE ensemble because, as was shown in Table 1, conformal methods have well controlled empirical coverage. The CE ensemble is drawn uniformly from the $1-\alpha$ confidence region (Section 3.3), so it is not simply concentrated near the median. In fact, the results in Table 1 suggest that the conformal prediction sets may even be slightly underconfident because the empirical sizes are generally greater than the nominal level. CE (Norm) shows slightly improved metrics compared with $\ell_{\infty}$-Depth and Tukey depth, but due to its consistent under-coverage towards the latter part of the prediction interval (2070-2100), this may be due to overconfidence.
@@ 4.3 Uncertainty quantification over time
As in Section 4.1, we break the disaggregate UQ skill results in Table 2 into decadal blocks to assess UQ skill over time. Because conformal methods can be susceptible to distribution shift, this will allow us to identify if there are any periods where CE performs worse than IMV or IMV(BC). We then further decompose our temporal results into monthly units, where SW is computed using only prediction sets and observations within a given month to assess seasonal variability. For CE, we use the $\ell_{\infty}$-Depth due to its overall more stable coverage (Figure 1) and comparable UQ skill compared to the other depths.
@@
The top row of Figure 2 shows the SW distance between the projection and the target over a given decade using each UQ method (IMV, IMV(BC), and CE) for each ensemble analysis function averaged over all model experiments. Although SW metrics tend to increase over time, particularly for TAS and TMAX projections, CE is consistently below IMV and IMV(BC) for all models and all time periods. Thus, even for far future projections (2070 - 2100), the climatologies produced by CE are much closer to the climatologies of the target process than if IMV had been used. One notable exception is LM using CE, which sits higher than all other models using CE in TAS and TMAX. This was likely caused by instabilities training LM, resulting in overly high or low projections at a small handful of spatial locations.
@@ 4.4 Marginal uncertainty quantification over space
Table 2 shows that, regardless of the ensemble analysis function or climate variable, swapping out IMV or IMV(BC) for a conformal ensemble (CE) uniformly decreases the approximation error of the projected climatology to the target processes climatology. Figure 2 confirms that this result holds for each decade in the projection interval (Apr. 2024 - Dec 2099). Because these results are based on the Sliced Wasserstein (SW) distance, they measure how close the projection’s joint distribution is to the target process’ joint distribution. They do not tell us how close their distributions are marginally at each grid location, i.e., if there are any spatially varied differences.